# Netan Dogra

### v2

@@ 1. Introduction
1.1. $2$-descent for hyperelliptic curves
@@
In contrast to the mysterious phenomena arising when considering ranks of abelian varieties, for Galois representations $M$ associated to pure motives of weight $<-1$, Bloch and Kato conjecture that the dimension of the Selmer group $H^{1}_{f}(\operatorname{Gal}(\overline{\mathbb{Q}}|\mathbb{Q}),M)$ is essentially a purely geometric invariant. More precisely, they conjecture a formula for it solely in terms of the action of complex conjugation on $M$, and the Hodge filtration on its de Rham realisation. There has been a vast amount of working on verifying this conjecture in various cases, which we will not attempt to survey here. However, both of the main approaches (via Euler systems, and via modularity lifting) depend on knowing that the underlying representation has an automorphic origin.
@@
where the outer vertical maps are the obvious ones. The main results on “$\Gamma$-descent” for $X$ proved in this paper may be viewed as nonabelian generalisations of the main results described above for $2$-descent on hyperelliptic curves. Firstly, we give an explicit field-theoretic description of the cohomology group $H^{1}(K,\wedge^{2}J[2])$, in terms of units mod squares of a certain étale algebra of degree $2g^{2}+g$ over $K$. If $\operatorname{Gal}(\overline{K}|K)$ acts $2$-transitively on the roots of the polynomial $f$, then this étale algebra is a field extension of $K$.
Given a pair of points $z,b$ in $X(K)$, we obtain a class $\kappa_{\Gamma,b}(z)$ in $H^{1}(K,\Gamma)$ whose image in $H^{1}(K,J[4])$ is the mod $4$ Kummer class $\kappa_{4,b}(z)$ of $[z]-[b]$. If $[z]-[b]$ lies in $4\cdot\mathrm{Jac}(X)(K)$, there is an element of $H^{1}(K,\wedge^{2}J[2])$ with image $\kappa_{\Gamma,b}(z)$ in $H^{1}(K,\Gamma)$. We give an explicit formula for such a class in the case when $X$ is hyperelliptic with a rational Weierstrass point, and the points $z$ and $b$ are not Weierstrass points.
@@
In the sequel to this paper [Dog23], we study liftability or ‘second descent’ obstructions for $H^{1}_{f}(\mathbb{Q}_{v},\wedge^{2}T_{2}J)$, and use these to verify finiteness of $X(\mathbb{Q}_{2})_{2}$ for several thousand genus 2 curves with Jacobians of Mordell–Weil rank 2 and Picard number 1.
@@ 2. Galois cohomology of the wedge square
This is given by sending $[x]\wedge[y]$ to [$[\{x,y\}]$. ∎
@@
It is enough to prove this when $S$ is a transitive $G$-set. In that case $S$ is isomorphic to $\operatorname{Gal}(K)\backslash\operatorname{Gal}(L)$ for a finite separable extension $L|K$, and $K_{S}$ is isomorphic to $L$ as above. On the other hand $\mathbb{F}_{2}[\operatorname{Gal}(K)\backslash\operatorname{Gal}(L)]$ is isomorphic to $\mathrm{Ind}^{L}_{K}\mathbb{F}_{2}$, and hence the statement follows from Shapiro’s lemma. ∎
@@
[…]
@@
[…]
@@
[…]
@@
(6) $0\to\wedge^{2}J[2]\to\wedge^{2}\mathrm{Ind}^{K_{f}}_{K}\mathbb{F}_{2}\to J[2]\to 0$
@@
3. The nonabelian $(x-T)$-map
In this section we give what may be viewed as a ‘nonabelian generalisation’ of the $(x-T)$ map defined by Schaefer [Sch95] generalising a construction of Cassels [Cas83].
@@ 3. The nonabelian $(x-T)$-map
For example, let $\Gamma$ denote the quotient defined in (1), and let $\kappa_{\Gamma,b}(z)$ in $H^{1}(K,\Gamma)$ denote the section class associated to $b$ and $z$. Then the image of $\kappa_{\Gamma,b}(z)$ in $H^{1}(K,J[4])$ is the mod $4$ Kummer class $\kappa_{4,b}(z)$ of $[z]-[b]$. If $[z]-[b]$ lies in $4\cdot\mathrm{Jac}(X)(K)$, there is an element of $H^{1}(K,\wedge^{2}J[2])$ with image $\kappa_{\Gamma,b}(z)$ in $H^{1}(K,\Gamma)$.
For any root $\gamma$ of $f$ over an étale $K$-algebra $L$, we let $c_{\gamma}:=x(b)-\gamma$. If $z\in X(K)$ is a non–Weierstrass point with the property that $z-b$ lies in $2\cdot J(K)$, then there is an element $u_{\gamma}$ of $L$ such that $c_{\gamma}u_{\gamma}^{2}=x(z)-\gamma$. For example, we can apply this construction to the étale algebra $K_{f,2}$, to get elements $c_{\alpha},u_{\alpha},c_{\beta}$ and $u_{\beta}$. The element $\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\beta-\alpha}$ will in fact be defined over the field $K_{f}^{(2)}$. This element of $K_{f}^{(2)}$ is related to $j_{J[4],b}(z)$ and $j_{\Gamma,b}(z)$ by the following ‘nonabelian $(x-T)$ maps’.
@@
[…]
@@
Isomorphism classes of $\mathbb{Z}/2\mathbb{Z}$ covers of $Y_{\alpha,K^{\mathrm{sep}}}$ unramified outside infinity are in bijection with equivalence classes of subsets of the set of roots of $\prod_{\beta\neq\alpha}(u_{\alpha}^{2}-(\beta-\alpha)/c_{\alpha})$, under the equivalence relation $S_{1}\sim S_{2}$ if and only if $\mathrm{Roots}(g_{\alpha})=S_{1}\sqcup S_{2}$. This bijection respects the action of $\mathrm{Aut}(Y_{\alpha}/X_{K_{f}})$ on both sets. The $\mathbb{Z}/2\mathbb{Z}$-covers of $Y_{\alpha,K^{\mathrm{sep}}}$ which come from a $\mathbb{Z}/4\mathbb{Z}$-cover of $X_{K^{\mathrm{sep}}}$ are exactly the $\mathrm{Aut}(Y_{\alpha}/X_{K_{f}})$-stable covers which do not descend to a $\mathbb{Z}/2\mathbb{Z}$-cover of $X_{K_{f}}$. The lemma follows from the fact that the generator of $\mathrm{Aut}(Y_{\alpha}/X_{K_{f}})$ acts on the set of roots of $\prod_{\beta\neq\alpha}(u_{\alpha}^{2}-(\beta-\alpha)/c_{\alpha})$ by swapping $(\frac{\beta-\alpha}{c_{\alpha}})^{1/2}$ and $-(\frac{\beta-\alpha}{c_{\alpha}})^{1/2}$ for all $\beta\neq\alpha$.
@@
[…]
@@
[…]
@@
Let $F$ be a free group on $2g$ generators $x_{1},\ldots,x_{2g}$. For each $i$, let $G_{i}$ denote the kernel of the homomorphism $F\to\mathbb{Z}/2\mathbb{Z}$ sending $x_{i}$ to $1$ and $x_{j}$ to 00 for $j\neq i$. Let $H_{i}<G_{i}$ denote the kernel of the mod 2 abelianisation map
@@
The kernel of $\Gamma\to F^{\operatorname{ab}}\otimes\mathbb{Z}/4\mathbb{Z}$ is an $\mathbb{F}_{2}$-vector space with basis $[x_{i},x_{j}]$ for $i\neq j$. On the other hand the image of $[x_{i},x_{j}]$ in $H_{i}^{\operatorname{ab}}\otimes\mathbb{F}_{2}$ is nonzero. Hence $H_{i}\supset\mathrm{Ker}(F\to\Gamma)$. ∎
@@ 3.1. Explicit formula for the $H^{1}(K,\wedge^{2}J[2])$ class
4. $2$-descent for Selmer schemes: local aspects away from $2$
@@ 4.1. Properties of $H^{1}_{f}(K_{v},T)$
5. $2$-descent for Selmer schemes: local aspects at $2$
To prove Theorem 1, it will be necessary to describe $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}J[2])$ as a subspace of $\mathbb{Q}_{2,f}^{(2),\times}\otimes\mathbb{F}_{2}$. Recall that by definition $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}J[2])$ is the image of $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}T_{2}J)$ in $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$. We aim to achieve this using the nonabelian $(x-T)$ map. The idea behind our approach is as follows. We can calculate the dimension of the image of $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}T_{2}J)$ in $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$, hence to describe the image it is enough to construct enough classes in $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$ which have crystalline lifts. We do this by using $\mathbb{Q}_{2}$ points of $X$ to construct crystalline classes in a larger Galois cohomology set, and try to give conditions for these to ‘induce’ crystalline classes with values in $\wedge^{2}J[2]$.
More precisely, let $b$ and $z$ be $\mathbb{Z}_{2}$ points of $X$ not reducing to the point at infinity, and let $\Pi_{2}$ denote the maximal $2$-unipotent quotient of $\pi_{1}^{\mathrm{\acute{e}t},2}(X_{\overline{\mathbb{Q}}_{2}}-\{\infty\},b)$. Then $j_{\Pi_{2},b}(z)$ is an element of $H^{1}_{f}(\mathbb{Q}_{2},\Pi_{2})$, i.e. an element of $H^{1}(\mathbb{Q}_{2},\Pi_{2})$ whose image in $H^{1}(\mathbb{Q}_{2},U_{2}(\mathbb{Q}_{2}))$ lies in the image of the subset $H^{1}_{f}(\mathbb{Q}_{2},U_{2}(\mathbb{Q}_{2}))$ in the sense of Kim [Kim09]. Furthermore, if $z-b$ lies in $4\cdot J(\mathbb{Q}_{2})$, then $j_{\Gamma,b}(z)$ lifts to an element of $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$, which we shall also denote $j_{\Gamma,b}(z)$ (although if $J[4](\mathbb{Q}_{2})\neq 0$ then this lift is not unique).
@@ 5.1. Integral mixed extensions
The issue with the idea outlined above is that the only way we have of constructing elements of $H^{1}_{f}(\mathbb{Q}_{2},\Pi_{2})$ is from $\mathbb{Z}_{2}$-points of the curve $X$. This means that the condition of finding a $\gamma$ with the same image in $H^{1}(\mathbb{Q}_{2},T_{2}J)$ as $[z]-[b]$ is too restrictive. For this reason, we replace $\Pi_{2}$ with a larger group for which it is easier to construct elements.
Let $\Lambda=\mathbb{F}_{p},\mathbb{Z}_{p}$ or $\mathbb{Q}_{p}$. Let $T_{0},\ldots,T_{n}$ be finite rank free $\Lambda$ module with a continuous action of a profinite group $G$. We define a mixed extension with graded pieces $T_{0},\ldots,T_{n}$ to a be a tuple $((M_{\bullet}),(\psi_{\bullet}))$ consisting of a finite free $\Lambda$ module with a continuous action of $G$ and a $G$ stable filtration
@@
together with $G$-equivariant isomorphisms of $\Lambda$-modules $\psi_{i}:M_{i-1}/M_{i}\stackrel{{\scriptstyle\simeq}}{{\longrightarrow}}T_{i}$ for all $1\leq i\leq n$. Given two mixed extensions $(M_{\bullet},\psi_{\bullet}),(M^{\prime}_{\bullet},\psi^{\prime}_{\bullet})$ with graded pieces $T_{0},\ldots,T_{n}$, we define a unipotent isomorphism between $M$ and $M^{\prime}$ to be an isomorphism of filtered $\Lambda$-modules $f:M_{0}\simeq M_{0}^{\prime}$ which is compatible with the $\psi_{i}$ and $\psi_{i}^{\prime}$. By compatible, we mean that, if $\mathrm{gr}_{i}f$ denotes the induced isomorphism $M_{i-1}/M_{i}\simeq M_{i-1}^{\prime}/M_{i}^{\prime}$, then we have $\psi_{i}=\psi_{i}^{\prime}\circ\mathrm{gr}_{i}f$ for all $i$.
@@
Now let $\Pi_{n}$ be the maximal pro-$p$ $n$-unipotent quotient of $\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{K}},b)$. Let $T_{i}$ denote the $i$th graded piece of the Iwasawa algebra $\mathbb{Z}_{p}[\![\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{X}}-\{\infty\},b)]\!]$ with respect to the $I$-adic filtration, where $I$ is the augmentation ideal. Define
@@
$1\to T_{0}^{*}\otimes T_{2}\to U(T_{\bullet})\to T_{0}^{*}\otimes T_{1}\oplus T_{1}^{*}\otimes T_{2}\to 1$
In terms of mixed extensions, a lift of an element in $H^{1}(G,T_{0}^{*}\otimes T_{1})\times H^{1}(G,T_{1}^{*}\otimes T_{2})$ to $H^{1}(G,U(T_{\bullet})$ corresponds to concatenating an extension $[E_{1}]$ of $T_{0}$ by $T_{1}$ and an extension $[E_{2}]$ of $T_{1}$ by $T_{2}$ to form a mixed extension with associated graded $T_{\bullet}$. Following Nekovář [Nek93] (following SGA [SGA72]), we shall refer to such a concatenation as a mixed extension of $E_{1}$ and $E_{2}$. The obstruction to forming such a mixed extension is exactly the image of $([E_{1}],[E_{2}])$ in $H^{2}(G,T_{0}^{*}\otimes T_{2})$ under the boundary map. We recall that this boundary map has the following simple description.
@@
We now turn to the problem of describing the structure of $E_{2}(b)$. By definition this is a direct sum of $\mathbb{Z}_{2}$ and $IE_{2}(b)$. Hence the problem is to describe the structure of $I\mathbb{Z}_{2}[\![\Pi_{2}]\!]/I^{3}$ as an extension of $T_{2}J$ by $T_{2}J^{\otimes 2}$. This is closely related to the class of the Ceresa cycle of $(X,b)$ in Galois cohomology, by work of Hain and Matsumoto [HM05]. The relation integrally (rather than over $\mathbb{Q}_{2}$) has some additional subtleties which were recently studied by Bisogno, Li, Litt and Srinivasan [BLLS23]. As we are working with hyperelliptic curves, the extension class is $2$-torsion, and we reduce to showing that this $2$-torsion element is in fact zero.
@@
Suppose $b\in X(K)$ is a Weierstrass point. Then $I\mathbb{Z}_{2}[\![\Pi_{2}]\!]/I^{3}$ splits.

### v3

# 22-descent for Bloch–Kato Selmer groups and rational points on hyperelliptic curves I

@@ 1. Introduction
$1.$ $2$-descent for hyperelliptic curves
@@
In contrast to the mysterious phenomena arising when considering ranks of abelian varieties, for Galois representations $M$ associated to pure motives of weight <1<-1, Bloch and Kato conjecture that the dimension of the Selmer group $H^{1}_{f}(\operatorname{Gal}(\overline{\mathbb{Q}}|\mathbb{Q}),M)$ is essentially a purely geometric invariant. More precisely, they conjecture a formula for it solely in terms of the action of complex conjugation on $M$, and the Hodge filtration on its de Rham realisation. There has been a vast amount of working on verifying this conjecture in various cases, which we will not attempt to survey here. However, both of the main approaches (via Euler systems, and via modularity lifting) depend on knowing that the underlying representation has an automorphic origin.
@@
where the outer vertical maps are the obvious ones. The main results on $\Gamma$-descent” for $X$ proved in this paper may be viewed as nonabelian generalisations of the main results described above for $2$-descent on hyperelliptic curves. Firstly, we give an explicit field-theoretic description of the cohomology group $H^{1}(K,\wedge^{2}J[2])$, in terms of units mod squares of a certain étale algebra of degree $2g^{2}+g$ over $K$. If $\operatorname{Gal}(\overline{K}|K)$ acts $2$-transitively on the roots of the polynomial $f$, then this étale algebra is a field extension of $K$.
Given a pair of points $z,b$ in $X(K)$, we obtain a class $\kappa_{\Gamma,b}(z)$ in $H^{1}(K,\Gamma)$ whose image in $H^{1}(K,J[4])$ is the mod $4$ Kummer class $\kappa_{4,b}(z)$ of [z][b][z]-[b]. If [z][b][z]-[b] lies in $4\cdot\mathrm{Jac}(X)(K)$, there is an element of $H^{1}(K,\wedge^{2}J[2])$ with image $\kappa_{\Gamma,b}(z)$ in $H^{1}(K,\Gamma)$. We give an explicit formula for such a class in the case when $X$ is hyperelliptic with a rational Weierstrass point, and the points $z$ and $b$ are not Weierstrass points.
@@
In the sequel to this paper [Dog23], we study liftability or ‘second descent’ obstructions for $H^{1}_{f}(\mathbb{Q}_{v},\wedge^{2}T_{2}J)$, and use these to verify finiteness of $X(\mathbb{Q}_{2})_{2}$ for several thousand genus 2 curves with Jacobians of Mordell–Weil rank 2 and Picard number 1. Recent work of Berry [Ber25] has shown how to extend many of these results to the case of hyperelliptic curves of higher genus, and of even degree (i.e. hyperelliptic curves without a rational Weierstrass point). Berry’s work also shows how the crystalline conditions, studied in section 5 using a nonabelian Kummer map, may be studied more systematically in the case where the curve has good ordinary reduction at 2.
@@ 2. Galois cohomology of the wedge square
This is given by sending $x]\wedge[y]$ to $\{x,y\}]$. ∎
@@
[…]
@@
$H^{1}(K,\mathbb{F}_{2}[S_{1}\times S_{2}])$$H^{1}(K,\mathbb{F}_{2}[S_{2}])$$(K_{S_{1}}\otimes K_{S_{2}})^{\times}\otimes\mathbb{F}_{2}$$K_{S_{2}}^{\times}\otimes\mathbb{F}_{2}.$}
@@
$H^{1}(K,\mathbb{F}_{2}[S^{(2)}])$$H^{1}(K_{2},\mathbb{F}_{2}[S])$$K_{f}^{(2),\times}\otimes\mathbb{F}_{2}$$K_{f}^{\times}\otimes\mathbb{F}_{2}.$}
@@
$H^{1}(K,\mathbb{F}_{2}[S^{(2)}])$$H^{1}(K,\mathbb{F}_{2}[S\times S])$$H^{1}(K_{2},\mathbb{F}_{2}[S])$$K_{f}^{(2),\times}\otimes\mathbb{F}_{2}$$(K_{f}\otimes K_{f})^{\times}\otimes\mathbb{F}_{2}$$K_{f}^{\times}\otimes\mathbb{F}_{2}.$}
@@
(6) $0\to\wedge^{2}J[2]\to\wedge^{2}\mathrm{Ind}^{K_{f}}_{K}\mathbb{F}_{2}\to J[2]\to 0.$
@@
3. The nonabelian (xT)(x-T)-map
In this section we give what may be viewed as a ‘nonabelian generalisation’ of the (xT)(x-T) map defined by Schaefer [Sch95] generalising a construction of Cassels [Cas83].
@@ 3. The nonabelian (xT)(x-T)-map
For example, let $\Gamma$ denote the quotient defined in (1), and let $\kappa_{\Gamma,b}(z)$ in $H^{1}(K,\Gamma)$ denote the section class associated to $b$ and $z$. Then the image of $\kappa_{\Gamma,b}(z)$ in $H^{1}(K,J[4])$ is the mod $4$ Kummer class $\kappa_{4,b}(z)$ of [z][b][z]-[b]. If [z][b][z]-[b] lies in $4\cdot\mathrm{Jac}(X)(K)$, there is an element of $H^{1}(K,\wedge^{2}J[2])$ with image $\kappa_{\Gamma,b}(z)$ in $H^{1}(K,\Gamma)$.
For any root $\gamma$ of $f$ over an étale $K$-algebra $L$, we let $c_{\gamma}:=x(b)-\gamma$. If $z\in X(K)$ is a non–Weierstrass point with the property that zbz-b lies in $2\cdot J(K)$, then there is an element $u_{\gamma}$ of $L$ such that $c_{\gamma}u_{\gamma}^{2}=x(z)-\gamma$. For example, we can apply this construction to the étale algebra $K_{f,2}$, to get elements $c_{\alpha},u_{\alpha},c_{\beta}$ and $u_{\beta}$. The element $\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\beta-\alpha}$ will in fact be defined over the field $K_{f}^{(2)}$. This element of $K_{f}^{(2)}$ is related to $j_{J[4],b}(z)$ and $j_{\Gamma,b}(z)$ by the following ‘nonabelian (xT)(x-T) maps’.
@@
$(\mathrm{Res}_{L|K}X)_{M}$$(\mathrm{Res}_{L|K}Y)_{M}$$\prod_{\sigma}X_{M}$$\prod_{\sigma}Y_{M}$$\scriptstyle{\simeq}$$\scriptstyle{(\mathrm{Res}_{L|K}f)_{M}}\scriptstyle{\simeq}$$\scriptstyle{\prod_{\sigma:L\to M}f_{\sigma}}$
@@
Isomorphism classes of $\mathbb{Z}/2\mathbb{Z}$ covers of $Y_{\alpha,K^{\mathrm{sep}}}$ unramified outside infinity are in bijection with equivalence classes of subsets of the set of roots of $\prod_{\beta\neq\alpha}(u_{\alpha}^{2}-(\beta-\alpha)/c_{\alpha})$, under the equivalence relation $S_{1}\sim S_{2}$ if and only if $\mathrm{Roots}(g_{\alpha})=S_{1}\sqcup S_{2}$. This bijection respects the action of $\mathrm{Aut}(Y_{\alpha}/X_{K_{f}})$ on both sets. The $\mathbb{Z}/2\mathbb{Z}$-covers of $Y_{\alpha,K^{\mathrm{sep}}}$ which come from a $\mathbb{Z}/4\mathbb{Z}$-cover of $X_{K^{\mathrm{sep}}}$ are exactly the $\mathrm{Aut}(Y_{\alpha}/X_{K_{f}})$-stable covers which do not descend to a $\mathbb{Z}/2\mathbb{Z}$-cover of $X_{K_{f}}$. The lemma follows from the fact that the generator of $\mathrm{Aut}(Y_{\alpha}/X_{K_{f}})$ acts on the set of roots of $\prod_{\beta\neq\alpha}(u_{\alpha}^{2}-(\beta-\alpha)/c_{\alpha})$ by swapping $(\frac{\beta-\alpha}{c_{\alpha}})^{1/2}$ and $(\frac{\beta-\alpha}{c_{\alpha}})^{1/2}$ for all $\beta\neq\alpha$.
@@
(11) $Y_{\alpha}$$\mathrm{Res}_{K_{f,2}|K_{f}}Y_{\alpha\beta,K_{f,2}}$$\mathbb{P}^{1}$$\mathrm{Res}_{K_{f,2}|K}\mathbb{P}^{1}.$$\scriptstyle{2(1+\frac{c_{\alpha}u_{\alpha}-c_{\beta}u_{\beta}}{\beta-\alpha})}$$\scriptstyle{\Delta}$
@@
(12) $X_{\Gamma}^{\prime}$$\mathrm{Res}_{K_{f,2}|K_{f}}Y_{\alpha\beta,K_{f,2}}$$X_{2}$$\mathrm{Res}_{K_{f,2}|K}X_{2,K_{f,2}}.$$\scriptstyle{\Delta}$
@@
Let $F$ be a free group on $2g$ generators $x_{1},\ldots,x_{2g}$. For each $i$, let $G_{i}$ denote the kernel of the homomorphism $F\to\mathbb{Z}/2\mathbb{Z}$ sending $x_{i}$ to $1$ and $x_{j}$ to 0 for $j\neq i$. Let $H_{i}<G_{i}$ denote the kernel of the mod 2 abelianisation map
@@
The kernel of $\Gamma\to F^{\operatorname{ab}}\otimes\mathbb{Z}/4\mathbb{Z}$ is an $\mathbb{F}_{2}$-vector space with basis $x_{i},x_{j}]$ for $i\neq j$. On the other hand the image of $x_{i},x_{j}]$ in $H_{i}^{\operatorname{ab}}\otimes\mathbb{F}_{2}$ is nonzero. Hence $H_{i}\supset\mathrm{Ker}(F\to\Gamma)$. ∎
@@ 3.1. Explicit formula for the $H^{1}(K,\wedge^{2}J[2])$ class
4. $2$-descent for Selmer schemes: local aspects away from 22
@@ 4.1. Properties of $H^{1}_{f}(K_{v},T)$
5. $2$-descent for Selmer schemes: local aspects at 22
To prove Theorem 1, it will be necessary to describe $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}J[2])$ as a subspace of $\mathbb{Q}_{2,f}^{(2),\times}\otimes\mathbb{F}_{2}$. Recall that by definition $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}J[2])$ is the image of $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}T_{2}J)$ in $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$. We aim to achieve this using the nonabelian (xT)(x-T) map. The idea behind our approach is as follows. We can calculate the dimension of the image of $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}T_{2}J)$ in $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$, hence to describe the image it is enough to construct enough classes in $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$ which have crystalline lifts. We do this by using $\mathbb{Q}_{2}$ points of $X$ to construct crystalline classes in a larger Galois cohomology set, and try to give conditions for these to ‘induce’ crystalline classes with values in $\wedge^{2}J[2]$.
More precisely, let $b$ and $z$ be $\mathbb{Z}_{2}$ points of $X$ not reducing to the point at infinity, and let $\Pi_{2}$ denote the maximal $2$-unipotent quotient of $\pi_{1}^{\mathrm{\acute{e}t},2}(X_{\overline{\mathbb{Q}}_{2}}-\{\infty\},b)$. Then $j_{\Pi_{2},b}(z)$ is an element of $H^{1}_{f}(\mathbb{Q}_{2},\Pi_{2})$, i.e. an element of $H^{1}(\mathbb{Q}_{2},\Pi_{2})$ whose image in $H^{1}(\mathbb{Q}_{2},U_{2}(\mathbb{Q}_{2}))$ lies in the image of the subset $H^{1}_{f}(\mathbb{Q}_{2},U_{2}(\mathbb{Q}_{2}))$ in the sense of Kim [Kim09]. Furthermore, if zbz-b lies in $4\cdot J(\mathbb{Q}_{2})$, then $j_{\Gamma,b}(z)$ lifts to an element of $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$, which we shall also denote $j_{\Gamma,b}(z)$ (although if $J[4](\mathbb{Q}_{2})\neq 0$ then this lift is not unique).
@@ 5.1. Integral mixed extensions
The issue with the idea outlined above is that the only way we have of constructing elements of $H^{1}_{f}(\mathbb{Q}_{2},\Pi_{2})$ is from $\mathbb{Z}_{2}$-points of the curve $X$. This means that the condition of finding a $\gamma$ with the same image in $H^{1}(\mathbb{Q}_{2},T_{2}J)$ as [z][b][z]-[b] is too restrictive. For this reason, we replace $\Pi_{2}$ with a larger group for which it is easier to construct elements.
Let $\Lambda=\mathbb{F}_{p},\mathbb{Z}_{p}$ or $\mathbb{Q}_{p}$. Let $T_{0},\ldots,T_{n}$ be finite rank free $\Lambda$ module with a continuous action of a profinite group $G$. We define a mixed extension with graded pieces $T_{0},\ldots,T_{n}$ to a be a tuple $((M_{\bullet}),(\psi_{\bullet}))$ consisting of a finite rank free $\Lambda$-module with a continuous action of $G$ and a $G$-stable filtration
@@
together with $G$-equivariant isomorphisms of $\Lambda$-modules $\psi_{i}:M_{i-1}/M_{i}\stackrel{{\scriptstyle\simeq}}{{\longrightarrow}}T_{i}$ for all $1\leq i\leq n$. Given two mixed extensions $(M_{\bullet},\psi_{\bullet}),(M^{\prime}_{\bullet},\psi^{\prime}_{\bullet})$ with graded pieces $T_{0},\ldots,T_{n}$, we define a unipotent isomorphism between $M$ and $M^{\prime}$ to be an isomorphism of filtered $\Lambda$-modules $f:M_{0}\simeq M_{0}^{\prime}$ which is compatible with the $\psi_{i}$ and $\psi_{i}^{\prime}$. By compatible we mean that, if $\mathrm{gr}_{i}f$ denotes the induced isomorphism $M_{i-1}/M_{i}\simeq M_{i-1}^{\prime}/M_{i}^{\prime}$, then we have $\psi_{i}=\psi_{i}^{\prime}\circ\mathrm{gr}_{i}f$ for all $i$.
@@
Now let $\Pi_{n}$ be the maximal pro-$p$ $n$-unipotent quotient of $\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{K}},b)$. Let $T_{i}$ denote the $i$th graded piece of the Iwasawa algebra $\mathbb{Z}_{p}[\![\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{K}}-\{\infty\},b)]\!]$ with respect to the $I$-adic filtration, where $I$ is the augmentation ideal. Define
@@
The image of the class of $\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{K}}-\{\infty\};b,z)$ is the class of the mixed extension
$E_{n}(b,z):=E_{n}(b)\times_{\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{K}}-\{\infty\},b)}\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{K}}-\{\infty\};b,z).$
@@
$1\to T_{0}^{*}\otimes T_{2}\to U(T_{\bullet})\to T_{0}^{*}\otimes T_{1}\oplus T_{1}^{*}\otimes T_{2}\to 1.$
In terms of mixed extensions, a lift of an element in $H^{1}(G,T_{0}^{*}\otimes T_{1})\times H^{1}(G,T_{1}^{*}\otimes T_{2})$ to $H^{1}(G,U(T_{\bullet})$ corresponds to concatenating an extension $E_{1}]$ of $T_{0}$ by $T_{1}$ and an extension $E_{2}]$ of $T_{1}$ by $T_{2}$ to form a mixed extension with associated graded $T_{\bullet}$. Following Nekovář [Nek93] (following SGA [SGA72]), we shall refer to such a concatenation as a mixed extension of $E_{1}$ and $E_{2}$. The obstruction to forming such a mixed extension is exactly the image of $([E_{1}],[E_{2}])$ in $H^{2}(G,T_{0}^{*}\otimes T_{2})$ under the boundary map. We recall that this boundary map has the following simple description.
@@
When working with mixed extensions over $\mathbb{Q}_{p}$, one can rescale the isomorphisms identifying the associated graded with $\oplus T_{i}$ to obtain a new mixed extension structure on the same representation. One new subtlety in the case of mixed extensions with $\mathbb{Z}_{p}$ rather than $\mathbb{Q}_{p}$ coefficients is that, given a mixed extension $M$, one can also construct new mixed extensions by choosing different lattices inside $M\otimes\mathbb{Q}_{p}$. This is used later to kill torsion. A simple case of this idea is presented in the following lemma.
@@
If $M$ is an integral mixed extension with graded pieces $T_{0},T_{1},T_{2}$, such that $M/T_{2}]\in p\cdot H^{1}_{f}(\mathbb{Q}_{p},T_{0}^{*}\otimes T_{1})$, then there is a Galois stable sublattice $M^{\prime}$ of $M$ with $M/M^{\prime}\simeq T_{1}\otimes\mathbb{F}_{p}$, and such that
$p[M^{\prime}/T_{2}]=[M/T_{2}]$
in $H^{1}_{f}(\mathbb{Q}_{p},T_{0}^{*}\otimes T_{1})$ and
$\mathrm{Ker}(M^{\prime}\to T_{0})]=p[\mathrm{Ker}(M\to T_{0})]$
in $H^{1}_{f}(\mathbb{Q}_{p},T_{1}^{*}\otimes T_{2})$.